OpenStudy (anonymous):

What is the equation for the line that passes through points (-1,-4) and (1, 4)? Write the equation in slope-intercept form! please help me!!

OpenStudy (anonymous):

Two point form of a line is y-y1 = ( y2-y1 / x2-x1) (x-x1) Substitute the given two points in the above equation. After substituting and getting the equation of line, change it into this form y = mx +c which is slope-intercept form

OpenStudy (anonymous):

64??